OpenStudy (anonymous):

Holly's city has 2.3 × 10^6 residents. Michael’s city has 3.5 × 10^5 fewer residents than Holly's city. How many people live in Michael’s town? A. 1.95 × 10^2 people B. 1.95 × 10^5 people C. 1.95 × 10^6 people <---- D. 1.95 × 10^7 people @shamim

OpenStudy (anonymous):

@jim_thompson5910 @ganeshie8 @ShadowLegendX @shamim @phi

OpenStudy (anonymous):

@mathslover @zepdrix @radar @kropot72 @tHe_FiZiCx99 @Lena772

OpenStudy (shamim):

2.3*10^6-3.5*10^5=?

OpenStudy (anonymous):

I think it's c.

OpenStudy (anonymous):

1950000

OpenStudy (anonymous):

@shamim

OpenStudy (anonymous):

\[\text{population of Michael’s town }=\text{population of Michael’s town }-3.5*10^5\]\[\text{population of Michael’s town }=2.3*10^6-3.5*10^5=(2.3-0.35)*10^6\]\[=1.95*10^6 \text{people}\]you are quite right

OpenStudy (shamim):

u r right

OpenStudy (anonymous):

A certain kind of passenger jet can travel at a speed of 8.4 × 10^5 meters in 1 hour. At this speed, how far can the jet travel in 5 hours? A. 4.2 × 10^4 meters B. 4.2 × 10^5 meters C. 4.2 × 10^6 meters D. 8.4 × 10^10 meters <----- @shamim

OpenStudy (anonymous):

@shamim @mukushla